### v1

@@ 6.2 Type I: control invariant set
This type of constraint restricts the state in a control invariant set, which is represented by the zero-sublevel set of feasibility function. There are mainly two kinds of feasibility functions of the control invariant set type: (a) control barrier function and (b) safety index. They differ in the inequality conditions that guarantee the invariant property of their zero-sublevel sets.
@@ 6.2.1 Control barrier function
Control barrier function (CBF) is a widely used feasibility function for synthesizing safe controllers in constrained OCPs (Ames et al., 2019, 2016; Cheng et al., 2019; Robey et al., 2020). The control invariance of the zero-sublevel set of CBF is guaranteed by the fact that the time derivative of CBF is always non-positive on the boundary of this set. As a state in the zero-sublevel set approaches the boundary, its CBF value may increase at first but will finally stop at some value not greater than zero. Therefore, the CBF will never take a value above zero as long as the initial state is in its zero-sublevel set. The formal definition of CBF in a discrete-time system is as follows.
Definition 6.3 (Control barrier function).
A function $B:\mathcal{X}\to\mathbb{R}$ is a control barrier function if it satisfies:
  1. 1.
@@
  2. 2.
    $\forall x\in\mathcal{X},\exists u\in\mathcal{U},\text{ s.t. }B(x^{\prime})-B(x)\leq-\alpha(B(x))$,
@@
If a function satisfies the above two properties, it can be proved that its zero-sublevel set is a control invariant set (Ames et al., 2019). Therefore, according to Definition 6.1, a CBF is a feasibility function itself, i.e.,
@@
Property 2) in Definition 6.3 is the most critical element for ensuring control invariance of the zero-sublevel set. This property tells us that there exists an action such that the increment of CBF from the current step to the next step is upper bounded. When solving a constrained OCP, we must take it as a virtual-time constraint to ensure that the obtained action actually satisfies this property. This leads to a one-step constraint imposed on the second state in virtual-time domain, i.e., $x_{1|t}$. In addition, $B(\cdot)\leq 0$ is imposed on the first state in virtual-time domain, i.e., $x_{0|t}$. This is because we must ensure that the first state is in the control invariant set so that the successive states can be kept in this set. This first-step virtual-time constraint is not weaker than the real-time state constraint because property 1) in Definition 6.3 tells us that $h(\cdot)>0$ implies $B(\cdot)>0$, which is equivalent to the statement that $B(\cdot)\leq 0$ implies $h(\cdot)\leq 0$. Combining these two constraints, we have
$g(x_{0|t})$ $=B(x_{0|t})\leq 0,$ (5)
@@
The above constraint guarantees that the next state of $x_{0|t}$, which is $x_{1|t}$, is still in the zero-sublevel set of CBF. If all states in the zero-sublevel set satisfy this constraint, the set is forward invariant, i.e., the state is always kept in this set. Therefore, the state constraint in real-time domain is always satisfied.
In literature, there is a notion of high order control barrier function (HOCBF) (Xiao and Belta, 2019) that is used for high relative degree constraints. Here, the relative degree of a function refers to its lowest order of time derivative where the action explicitly appears. When the constraint function $h$ has a high relative degree, directly choosing $h$ as a CBF cannot effectively constrain the action because constraint (5) is irrelevant to the action. This necessitates HOCBF, which considers time derivatives of $h$ where the action explicitly appears. When the time derivatives satisfy certain properties that ensure control invariance, $h$ is an HOCBF (Xiao and Belta, 2019). Although not explicitly stated, Definition 6.3 is actually compatible with HOCBF because $B$ does not have to be chosen as $h$ and can be constructed as some function that makes the action explicitly appear in constraint (5).
Let us analyze the size of IFR and EFR under constraint (5). For IFR, we look for the state that if we take it as the starting point of virtual-time OCP with constraint (5), this OCP has a solution. In the first step, all states in the zero-sublevel set of $B$ satisfy the virtual-time constraint. For notation simplicity, we denote this set as $X_{B}=\{x\in\mathcal{X}|B(x)\leq 0\}$. In the second step, according to property 2) in Definition 6.3, all states in the entire state space satisfy the virtual-time constraint. The IFR is the intersection of these two sets, which is $\mathrm{X}_{B}$, i.e.,
@@
For EFR, we look for the state from which the OCP with constraint (5) has a solution in all successive states as long as an initially feasible policy is applied in every step. According to Theorem 5.1, EFR must be a subset of IFR. Since the IFR under constraint (5) is $\mathrm{X}_{B}$, the EFR must be a subset of $\mathrm{X}_{B}$. For any state in $\mathrm{X}_{B}$, its successive state is still in $\mathrm{X}_{B}$ if we take an action such that the second constraint in (5) is satisfied. This requirement for the action is satisfied by any initially feasible policy. Thus, the state can always be kept in $\mathrm{X}_{B}$ as long as an initially feasible policy is applied at every time step. Moreover, since $\mathrm{X}_{B}$ is the IFR, the OCP with constraint (5) always has a solution in it. Therefore, we can conclude that the EFR equals $\mathrm{X}_{B}$, i.e.,
@@
We can see that EFR equals IFR under constraint (5). This shows the benefit of feasibility function for designing virtual-time constraints, i.e., it ensures the equivalence of EFR and IFR. This equivalence enables us to know the size of EFR by examining that of IFR. In a constrained OCP, we always aim to construct an EFR that equals the maximum EFR so that the policy obtained by solving the OCP has the largest working area. However, the size of EFR is difficult to obtain because it involves checking the constraint satisfaction in an infinite horizon. With the equivalence of EFR and IFR, we can avoid this difficulty by obtaining the size of IFR instead. If we construct an IFR that equals the maximum EFR, then the equivalence of EFR and the maximum EFR is also ensured. For a virtual-time constraint designed by CBF, its IFR is the zero-sublevel set of the CBF, i.e., $\mathrm{X}_{B}$. The size of this set depends on the design of CBF. A better design results in a larger $\mathrm{X}_{B}$, which is closer to the maximum EFR. In the best case, $\mathrm{X}_{B}$ equals the maximum EFR. For a general CBF, we can only say that $\mathrm{X}_{B}$ is a subset of the maximum EFR, i.e.,
@@
In conclusion, a CBF is a feasibility function that represents EFR using a control invariant set, which is its zero-sublevel set. A CBF must satisfy two properties, of which the first requires that the states in the control invariant set are constraint-satisfying, and the second requires that the increase rate of CBF is upper bounded under some action so that it will never exceed zero in the control invariant set. With a CBF, we can construct a two-step virtual-time constraint that restricts the state in the control invariant set in the first step and requires the action to satisfy the increase rate bound in the second step. Both IFR and EFR of this virtual-time constraint equal the control invariant set of CBF. However, this set may not equal the maximum EFR, i.e., it may be smaller than the maximum EFR. The size of control invariant set depends on the design of CBF. In systems with high-dimensional state and action spaces or systems without analytical models, it is usually difficult to handcraft a CBF whose control invariant set is equal to or close to the maximum EFR. Designing a CBF with the maximum control invariant set, i.e., the maximum EFR, in these systems remains a challenge.
@@ 6.2.2 Safety index
Safety index, originally proposed by Liu et al. (Liu and Tomizuka, 2014), is another feasibility function that represents a control invariant set by its zero-sublevel set, which is called safe set. Safety index ensures control invariance in a similar way to CBF. In particular, it is required that once the system deviates from the safe set, the control policy will pull it back. This is achieved by setting the time derivative of safety index to be negative outside the safe set. Safety index differs from CBF in that it explicitly considers the relative degree of system in construction. For systems with a high relative degree, the first-order time derivative is not enough to ensure control invariance because it does not constrain the action. To deal with this issue, safety index adopts a specific function form, which contains the constraint function or its nonlinear substitution, and a linear combination of its time derivatives. The formal definition of safety index is as follows.
Definition 6.4 (safety index).
A function $\phi:\mathrm{X}\to\mathbb{R}$ is a safety index if it satisfies the following three properties:
  1. 1.
@@
  2. 2.
    All roots of $01+k_{1}s+k_{2}s^{2}+\dots+k_{n}s^{n}=$ 0 are on the negative real line,
  3. 3.
@@
wher $n+1$ is the system order. Moreover, suppose $h^{*}$ defines the same set as $h$ does, i.e., $\{x\in\mathcal{X}|h^{*}(x)\leq 0\}=\mathrm{X}_{\mathrm{cstr}}$, then
@@
is also a safety index.
Property 3) above ensures that the relative degree of safety index is one so that the action can be constrained by its first-order time derivative. With the above three properties, it can be proved that when the action $u$ is unconstrained, there always exists $u$ that satisfies $\dot{\phi}\leq 0$ when $\phi=0$ (Liu and Tomizuka, 2014). When $u$ is constrained in some space $\mathcal{U}$, a 4-th condition needs to be added: the parameters $k_{i},i=1,2,\dots,n$ are chosen such that there exist $u\in\mathcal{U}$ for $\dot{\phi}\leq 0$ when $\phi=0$. This is a necessary and sufficient condition for control invariance of zero-sublevel set of $\phi$, i.e., the safe set. In Definition 6.4, the use of $h^{*}$ is to introduce nonlinearity to the safety index. Since it defines the same set as $h$, it only shapes the boundary of the feasible region. With a safety index, we can construct the following feasibility function:
@@
Here, including $h$ in the maximum operator is to ensure that the feasible region is a subset of the constrained set. When $h^{*}=h$ (i.e. the safety index is linearly defined), the feasible region of safety index can be represented by a set of inequalities similar to that of HOCBF (Xiao and Belta, 2019). When $h^{*}\neq h$ (i.e., the safety index is nonlinearly defined) and $n=1$ (i.e., the dynamic system is second order), the feasible region of safety index equals the zero-sublevel set of $F$. For $h^{*}\neq h$ and $n>1$, there is no conclusion regarding the exact representation of the feasible region with $F$.
The control invariance of safe set is guaranteed by the inequality $\dot{\phi}\leq 0$ when $\phi=0$. This is a differential inequality and only applies to continuous-time systems. To make it applicable to discrete-time systems, we must convert it to a difference inequality. An intuitive conversion is to replace the derivative of $\phi$ with its discretization obtained by, for example, the forward Euler method, which results in the following inequality, $\phi(x_{1|t})-\phi(x_{0|t})\leq 0$ when $\phi(x_{0|t})=0$. However, this inequality only constrains the value of $\phi$ when $\phi(x_{0|t})$ is exactly zero, which is not enough to guarantee the control invariance of safe set in discrete-time systems. This is because even if $\phi(x_{0|t})<0$, it is possible that $\phi$ becomes positive after only one step at $x_{1|t}$. Therefore, we must also constrain the value of $\phi$ when $\phi(x_{0|t})<0$. Specifically, we require that $\phi(x_{1|t})<0$ so that $x_{1|t}$ is still in the safe set. When $\phi(x_{0|t})>0$, $x_{0|t}$ is already out of the safe set. In this case, $\phi$ should decrease fast enough so that the state can return to the safe set quickly. Specifically, we require that $\phi(x_{1|t})\leq\phi(x_{0|t})-\eta$, where $\eta$ is a positive real number. This inequality can be further relaxed to $\phi(x_{1|t})\leq\max\left\{\phi(x_{0|t})-\eta,0\right\}$ because once $\phi(x_{1|t})\leq 0$, the state already enters the safe set. The above two inequalities can be written as the following inequality,
@@
which acts as a virtual-time constraint of safety index. Similar to the virtual-time constraint of CBF, the safety index also needs a first-step constraint that restricts the state in the safe set. Combining these two constraints, we arrive at the virtual-time constraint of safety index,
$g(x_{0|t})$ $=\max\{\phi(x_{0|t}),h(x_{0|t})\}\leq 0,$ (6)
@@
The above constraint satisfies the requirement that the virtual-time constraint in the first time step must not be weaker than the real-time constraint. The second constraint in (6) requires that the safety index decreases every time step by at least a value of $\eta$ when it is above zero. When the safety index is below zero, it is required not to become positive. With a policy satisfying constraint (6), the zero-sublevel set of safety index is forward invariant.
The relationship between feasible regions under safety index constraint is similar to that of CBF, so we omit the analysis here. One may discover that safety index is closely related to CBF in both feasibility function formulation and virtual-time constraint design. This relationship is not a coincidence but rather originates from the challenge of constructing a CBF. While CBF guarantees control invariance of its zero-sublevel set by definition, it does not provide any guidance for its construction. HOCBF provides one design rule for obtaining a valid CBF. Safety index can be considered as another approach to achieving this objective. However, we have seen that these methods, which define feasibility functions through control invariant sets, may not be able to obtain the maximum EFR. In contrast, the following feasibility function defined by constraint aggregation can lead to the maximum EFR.
@@ 6.3 Type II: constraint aggregation
This type of constraint replaces the infinite-step real-time constraints with a single-step virtual-time constraint by an aggregation function. Under this perspective, there are mainly two kinds of feasibility functions of this type: cost value function and Hamilton-Jacobi reachability function. The former uses a discounted summation function as the aggregation function while the latter uses a maximum function as the aggregation function. There are two types of virtual-time constraints constructed by this kind of feasibility function. One type uses the feasibility function of the current policy, which is to be optimized, to construct constraints. The other uses the feasibility function of a fixed policy, e.g., the policy from the last iteration, to construct constraints.
@@ 6.3.1 Cost value function
Cost value function (CVF) originates from a common formulation for safe RL, constrained Markov decision process (CMDP) (Altman, 1999), which augments a standard MDP with a cost function in parallel with the reward function. Being the expected cumulative costs, CVF shares the same mathematical form, and therefore the same properties, with state-value function $V^{\pi}(x)$, which greatly facilitates its usage. With a slight effort to define a cost function $c(x)$ with the constraint function $h(x)$, CVF can be well adapted to constrained OCPs. The definition of CVF is as follows.
Definition 6.5 (cost value function).
@@
To see that CVF is a valid feasibility function, note that $c(x)$ is non-negative for all $x$, so that $F^{\pi}(x)\leq 0$ is sufficient and necessary for $c(x_{t})=0,\forall t=0,1,\dots,\infty$. The close relationship between CVF and state-value function implies some common properties. Actually, it is indeed the fact that self-consistency condition and Bellman equation, which play important roles in solving state-value functions, have their respective counterparts when it comes to CVFs. For the sake of distinction, they are termed risky self-consistency condition:
@@
Similar to the case for state-value function, the right-hand sides of the above two equations can also be viewed as contraction operators on a complete metric space. Hence, by fixed-point iteration, a CVF can be solved just like a state-value function.
Until now, we have defined a valid feasibility function with CVF and showed that it is practically available, the next thing is to further construct a virtual-time constraint with it. The first way of constructing a virtual-time constraint is
$g(x_{0|t})=F^{\pi}(x_{0|t})\leq 0,$ (7)
and $x_{i|t},i=1,\dots,n/\infty$ are unconstrained. Here, the superscript $\pi$ of $F$ is the policy to be optimized. Although $x_{0|t}$ is not affected by $\pi$, constraint (7) is related to $\pi$ because $F^{\pi}$ is a function of $\pi$. Technically speaking, for a given $x_{0|t}$, $F^{\pi}(x_{0|t})$ is a functional of $\pi$. In practice, $F^{\pi}$ is difficult to compute directly and is usually approximated by importance sampling of trajectories collected by another policy, similar to the way of approximating value function in on-policy RL algorithms. This form of virtual-time constraint is compatible with direct RL/ADP methods for constrained OCPs so that it can serve as an equivalent substitute for the real-time constraint. Through the aggregation of constraints introduced by the summation function, the horizon of virtual-time constraint reduces to only one step, the current step. The IFR under constraint (7) is by definition simply
@@
And no matter which initially feasible policy we apply at the current step, we will always reach a state in the IFR. This is because this policy must still be initially feasible in the next state, which is guaranteed by the fact that the CVF’s being less than or equal to 00 at a single state renders all future states safe. As a result, the EFR is still
@@
What’s more, an excellent property of constraint (7) is that
@@
The first equation is obvious from the above analysis. To see why the second equation holds, recall Theorem 4.1 and that $F^{\pi}(x)\leq 0$ is equivalent to $h(x_{t})\leq 0,t=0,1,\dots,\infty$. The equivalence of these three regions is satisfying, for we can easily check the endless feasibility of a state by its initial feasibility, and the EFR is the maximum one that can ever be obtained.
The second way of constructing a virtual-time constraint is a little bit complex, involving both the current step and the next step:
$g(x_{i|t})=F^{k}(x_{i|t})\leq 0,i=0,1,$ (8)
and $x_{i|t},i=2,\dots,n/\infty$ are unconstrained. Unlike the former type, here the constraint function, which is the CVF of a fixed policy $\pi_{k}$, is irrelevant to the policy $\pi$ to be optimized. Although $F^{k}(x_{0|t})$ is actually unrelated with $\pi$, $x_{1|t}$ is induced by $\pi$, which makes constraint (8) effective. Under this constraint, for a state $x$ to be initially feasible, the first requirement is that $g(x_{0|t})=F^{k}(x_{0|t})\leq 0$, where $x_{0|t}=x$. This implies that if we take $\pi=\pi_{k}$, we will have $F^{k}(x_{1|t})\leq 0$. Consequently, any $x$ satisfying $F^{k}(x)\leq 0$ is an initially feasible state, with $\pi_{k}$ being an initially feasible policy in it. That is,
@@
Unfortunately, under this situation, we don’t have as good equivalence to the maximum EFR as above. The IFR and EFR are related to a fixed policy which can be highly unsatisfying, rendering only a few states safe. So, we can only conclude that
@@
However, by properly choosing $\pi_{k+1}$ under the guide of $F^{\pi_{k}}(x)$, the fact is that we can ensure the monotonic expansion of EFR and it will converge to the maximum EFR (Yang et al., 2023b).

### v2

@@ 6.2 Type I: control invariant set
This type of constraint restricts the state in a control invariant set (CIS) represented by the zero-sublevel set of a feasibility function. There are mainly two kinds of feasibility functions of this type: a) CBF and b) SI. They differ in the inequalities that guarantee the invariant property of their zero-sublevel sets.
@@ 6.2.1 Control barrier function
CBF is a widely used feasibility function for synthesizing safe controllers in constrained OCPs [ames2019control, ames2016control, cheng2019end, robey2020learning]. The control invariance of the zero-sublevel set of CBF is guaranteed by the fact that the time derivative of CBF is always non-positive on the boundary of this set. The formal definition of CBF in a discrete-time system is as follows.
Definition 6.2.1 (Control barrier function).
A function $B:\mathcal{X}\to\mathbb{R}$ is a control barrier function (CBF) if it satisfies:
  1. 1)
@@
  2. 2)
[…]
@@
If a function satisfies the above two properties, it can be proved that its zero-sublevel set is a CIS [ames2019control]. Therefore, according to Definition 6.1.1, a CBF is a feasibility function itself, i.e.,
@@
Property 2) in Definition 6.2.1 is the most critical element for ensuring control invariance. When solving a constrained OCP, we must take it as a virtual-time constraint to ensure that the action satisfies this property. This leads to a one-step constraint imposed on the second state in virtual-time domain, i.e., $x_{1|t}$. In addition, $B(\cdot)\leq 0$ is imposed on the first state $x_{0|t}$. This is because we need to restrict the domain of the OCP to the CIS. This first-step virtual-time constraint is not weaker than the real-time state constraint because property 1) in Definition 6.2.1 tells us that $B(\cdot)\leq 0$ implies $h(\cdot)\leq 0$. Combining these two constraints, we have
$g(x_{0|t})$ $=B(x_{0|t})\leq 0,$ (6.1)
@@
The above constraints guarantee that the next state $x_{1|t}$ is still in the zero-sublevel set. If all states in the zero-sublevel set satisfy this constraint, the set is forward invariant, and the real-time state constraint is always satisfied.
In literature, a notion of high order CBF (HOCBF) [xiao2019control, xiao2021high] is used to deal with high relative degree constraints. The relative degree of a function refers to its lowest order of time derivative where the action explicitly appears. When the constraint function $h$ has a high relative degree, directly choosing $h$ as a CBF has no effect on the action because constraint (6.1) is irrelevant to the action. This necessitates HOCBF, which considers time derivatives of $h$ where the action explicitly appears [xiao2019control]. Although not explicitly stated, Definition 6.2.1 is compatible with HOCBF because $B$ does not have to be chosen as $h$ and can be constructed as some function that makes the action explicitly appear in constraint (6.1).
Let us analyze the IFR and EFR under constraint (6.1). For IFR, we examine each state $x$ whether the corresponding virtual-time OCP starting from it has a solution. Regarding the first inequality in (6.1), since $x_{0|t}=x$, all states in the zero-sublevel set of $B$ naturally satisfy this condition. For notation simplicity, we denote this set as $\mathrm{X}_{B}=\{x\in\mathcal{X}|B(x)\leq 0\}$. Regarding the second inequality in (6.1), Definition 6.2.1 guarantees that for every state $x_{0|t}=x\in\mathcal{X}$, there exists an action such that the resulting state $x_{1|t}$ satisfies this inequality. Thus, we can conclude that
@@
For EFR, we look for the state where the OCPs in all successive states have solutions. Since the IFR is $\mathrm{X}_{B}$, the EFR must be a subset of $\mathrm{X}_{B}$. For any state in $\mathrm{X}_{B}$, its successive state is still in $\mathrm{X}_{B}$ if we take an action that satisfies (6.1). Therefore, the state can always be kept in $\mathrm{X}_{B}$ as long as an initially feasible policy is applied at every time step. We can conclude that the EFR equals $\mathrm{X}_{B}$, i.e.,
@@
We can see that EFR equals IFR under constraint (6.1). This shows the benefit of feasibility function for designing virtual-time constraints, i.e., it ensures the equivalence of EFR and IFR. In a constrained OCP, we always aim to construct an EFR that equals the maximum EFR. With the equivalence of EFR and IFR, we can achieve this by constructing an IFR that equals the maximum EFR. The size of IFR depends on the design of CBF. For a general CBF, we can only say that $\mathrm{X}_{B}$ is a subset of the maximum EFR, i.e.,
@@
In conclusion, a CBF is a feasibility function that represents EFR using a CIS, which is its zero-sublevel set. A CBF satisfies two properties: the first requires that the zero-sublevel set is constraint-satisfying, and the second ensures that the zero-sublevel set is control invariant. With a CBF, we can construct a two-step virtual-time constraint under which the IFR and EFR both equal the zero-sublevel set. However, this set may not equal the maximum EFR but only be a subset of it depending on the design of CBF. In systems with high-dimensional state and action spaces or systems without analytical models, it remains challenging to design a CBF with a CIS equal to or close to the maximum EFR.
@@ 6.2.2 Safety index
SI, originally proposed by liu2014control, is another feasibility function that represents a CIS by its zero-sublevel set, which is called safe set. SI ensures control invariance in a similar way to CBF. In particular, it requires that once the system deviates from the safe set, the control policy will pull it back. This is achieved by setting the time derivative of SI to be negative outside the safe set. SI differs from CBF in that it explicitly considers the relative degree of system in construction. It is constructed as the linear combination of the constraint function, its nonlinear substitution, and its time derivatives. The formal definition of SI is as follows.
Definition 6.2.2 (Safety index).
A function $\phi:\mathrm{X}\to\mathbb{R}$ is a safety index (SI) if it satisfies the following three properties:
  1. 1)
@@
  2. 2)
    All roots of $1+k_{1}s+k_{2}s^{2}+\dots+k_{n}s^{n}=0$ are on the negative real line,
  3. 3)
@@
where $n+1$ is the system order. Moreover, suppose $h^{*}$ defines the same set as $h$ does, i.e., $\{x\in\mathcal{X}|h^{*}(x)\leq 0\}=\mathrm{X}_{\mathrm{cstr}}$, then
@@
is also an SI.
Property 3) ensures the SI has a relative degree of one. This means that the action has a direct influence on the SI. With the above three properties, it can be proved that for unbounded action space, the safe set is control invariant [liu2014control]. For bounded action space $\mathcal{U}$, a fourth condition is needed: the parameters $k_{i},i=1,2,\dots,n$ are chosen such that there exist $u\in\mathcal{U}$ for $\dot{\phi}\leq 0$ when $\phi=0$, which is a necessary and sufficient condition for control invariance of the safe set. The use of $h^{*}$ is to introduce nonlinearity for shaping the boundary of the feasible region. With an SI, we can construct the following feasibility function:
@@
Here, including $h$ in the maximum operator is to ensure that the feasible region is a subset of the constrained set. When $h^{*}=h$ (i.e., the SI is linearly defined), the feasible region can be represented by a set of inequalities similar to that of HOCBF [xiao2019control]. When $h^{*}\neq h$ (i.e., the SI is nonlinearly defined) and $n=1$ (i.e., the dynamic system is second order), the feasible region equals the zero-sublevel set of $F$. For $h^{*}\neq h$ and $n>1$, there is no conclusion regarding the exact representation of the feasible region with $F$.
The control invariance of the safe set is guaranteed by the inequality $\dot{\phi}\leq 0$ when $\phi=0$. This is a differential inequality and only applies to continuous-time systems. For discrete-time systems, we can convert it to a difference inequality through a forward Euler discretization:
@@
Similar to the virtual-time constraint of CBF, SI also needs a first-step constraint that restricts the state in the safe set. Combining these two constraints, we arrive at the virtual-time constraint of SI,
$g(x_{0|t})$ $=\max\{\phi(x_{0|t}),h(x_{0|t})\}\leq 0,$ (6.2)
@@
The first constraint above is not weaker than the real-time constraint. The second constraint requires that the SI decreases every time step by at least a value of $\eta$ when it is above zero. When the SI is below zero, it is required not to become positive. With a policy satisfying constraint (6.2), the zero-sublevel set of SI is forward invariant.
The relationship between feasible regions under SI constraint is similar to that of CBF, so we omit the analysis here. One may discover that SI is closely related to CBF in both feasibility function formulation and virtual-time constraint design. This is not a coincidence but rather originates from the construction of CBFs. While CBF guarantees control invariance of its zero-sublevel set by definition, it does not provide any guidance for its construction. HOCBF provides one design rule for obtaining a valid CBF, and SI can be considered as another. However, we have seen that a limitation of these CIS-based methods is that they may not be able to obtain the maximum EFR. In contrast, the following feasibility function defined by constraint aggregation can lead to the maximum EFR.
@@ 6.3 Type II: constraint aggregation
Constraints based on constraint aggregation (CA) replace the infinite-step real-time constraints with a single-step virtual-time constraint by an aggregation function. There are mainly three kinds of CA-based feasibility functions: cost value function, Hamilton-Jacobi reachability function, and constraint decay function. Each kind of feasibility function further corresponds to two types of virtual-time constraints. One uses the feasibility function of the current policy, i.e., the policy to be optimized, while the other uses the feasibility function of a fixed policy, e.g., the policy from the last iteration.
An advantage of this type of feasibility functions is that they provide a principled way to evaluate the feasible region of a given policy without a known system model. These feasibility functions are typically associated with a self-consistency condition that depends only on the policy to evaluate and on sampled transitions from the environment. This equation induces a contraction mapping, so its fixed point can be learned via standard RL temporal-difference methods without requiring a system model [fisac2019bridging, yang2023feasible]. This enables us to solve these feasibility functions by fixed point iteration, Once the feasibility function is solved, its zero-sublevel set represents the feasible region of the policy.
@@ 6.3.1 Cost value function
Cost value function (CVF) originates from a common formulation for safe RL, constrained Markov decision process (CMDP) [altman1999constrained], which augments a standard MDP with a cost function. Being the expected cumulative costs, CVF shares the same mathematical form and, therefore, the same properties, with state-value function $V^{\pi}(x)$ in RL. With a slight effort to define a cost function $c(x)$ with the constraint function $h(x)$, CVF can be well adapted to constrained OCPs. The definition of CVF is as follows.
Definition 6.3.1 (Cost value function).
@@
To see that CVF is a valid feasibility function, note that $c(x)$ is non-negative for all $x$, so that $F^{\pi}(x)\leq 0$ is sufficient and necessary for $c(x_{t})=0,\forall t=0,1,\dots,\infty$. The close relationship between CVF and state-value function implies some common properties. Specifically, the self-consistency condition and Bellman equation, which play important roles in solving state-value functions, have their counterparts for CVFs. For the sake of distinction, they are termed risky self-consistency condition:
@@
Similar to the case for state-value function, the right-hand sides of the above two equations are contraction operators on a complete metric space. Hence, a CVF can be solved by fixed-point iteration, just like a state-value function.
There are two ways to construct virtual-time constraints with a CVF. The first way is to constrain only the first state:
$g(x_{0|t})=F^{\pi}(x_{0|t})\leq 0,$ (6.3)
and $x_{i|t},i=1,\dots,n$ are unconstrained. Here, the superscript $\pi$ of $F$ is the policy to be optimized. The dependence of constraint (6.3) on $\pi$ is established through $F^{\pi}$, which is a function of $\pi$. Technically speaking, for a given $x_{0|t}$, $F^{\pi}(x_{0|t})$ is a functional of $\pi$. In practice, $F^{\pi}$ is difficult to compute directly and is usually approximated by importance sampling of trajectories collected by another policy, similar to the way of approximating value function in on-policy RL algorithms. The IFR under constraint (6.3) is by definition
@@
Starting from $\mathrm{X}_{\mathrm{init}}^{g}$, as long as we apply an initially feasible policy, the successive states will always be in $\mathrm{X}_{\mathrm{init}}^{g}$. This is because $F^{\pi}(x)\leq 0$ is already an infinite-horizon constraint-satisfying property and naturally holds for all future steps. Therefore, the EFR is still
@@
What’s more, an excellent property of constraint (6.3) is that
@@
To understand this, recall Theorem 4.2.4 and that $F^{\pi}(x)\leq 0$ is equivalent to $h(x_{t})\leq 0,t=0,1,\dots,\infty$. The equivalence of these three regions is satisfying, with which we can easily check endless feasibility by initial feasibility, and a constructed EFR is automatically the maximum.
The second way of constructing a virtual-time constraint is a little bit complex, involving both the first and the second states:
$g(x_{i|t})=F^{k}(x_{i|t})\leq 0,i=0,1,$ (6.4)
and $x_{i|t},i=2,\dots,n$ are unconstrained. Unlike the former way, here $F^{k}$ is the CVF of a fixed policy $\pi_{k}$, which is irrelevant to the policy $\pi$ to be optimized. The dependence of constraint (6.4) on $\pi$ is established through $x_{1|t}$, which is induced by $\pi$. Under this constraint, any $x$ satisfying $F^{k}(x)\leq 0$ is initially feasible, with $\pi_{k}$ being an initially feasible policy. Therefore,
@@
Unfortunately, under constraint (6.4), the IFR and EFR depend on the policy $\pi_{k}$ and may be only subsets of the maximum EFR, i.e.,
@@
However, by properly choosing $\pi_{k+1}$ under the guide of $F^{\pi_{k}}(x)$, we can achieve a sequence of monotonically expanding EFRs, which converge to the maximum EFR [yang2023feasible].